how do you simplify 36.2 trillion/98.5 billion then write it in scientific notation?
Found 2 solutions by checkley75, Smirnov:
Answer by checkley75(3666)

36.2^10^15/98.5*10^12
36.2/98.5=.3675126 &
10^15/10^12=10^3
.3675126*10^3=367.5126= 3.675126*10^2

Answer by Smirnov(15)

36,200,000,000,000/98,500,000,000
Would be 72400/197.